Henry's sister is 3 years younger than Henry. The product of their ages is 180. How old is Henry?
help solve this or equation set up.
h(h+3)= 180

Answer by stanbon(75887)

h(h+3)= 180
h^2+3h-180 = 0
(h-12)(h+15) = 0
---
Positive solution:
h = 12 yrs old
h+3 = 15 yrs old
======================
Cheers,
Stan H.